# Marvel: Accelerating Safe Online Reinforcement Learning with Finetuned Offline Policy

### v2

@@ Abstract
The high costs and risks involved in extensive environment interactions hinder the practical application of current online safe reinforcement learning (RL) methods. While offline safe RL addresses this by learning policies from static datasets, the performance therein is usually limited due to reliance on data quality and challenges with out-of-distribution (OOD) actions. Inspired by recent successes in offline-to-online (O2O) RL, it is crucial to explore whether offline safe RL can be leveraged to facilitate faster and safer online policy learning, a direction that has yet to be fully investigated. To fill this gap, we first demonstrate that naively applying existing O2O algorithms from standard RL would not work well in the safe RL setting due to two unique challenges: erroneous Q-estimations, resulted from offline-online objective mismatch and offline cost sparsity, and Lagrangian mismatch, resulted from difficulties in aligning Lagrange multipliers between offline and online policies. To address these challenges, we introduce Marvel, a novel framework for O2O safe RL, comprising two key components that work in concert: Value Pre-Alignment to align the Q-functions with the underlying truth before online learning, and Adaptive PID Control to effectively adjust the Lagrange multipliers during online finetuning. Extensive experiments demonstrate that Marvel significantly outperforms existing baselines in both reward maximization and safety constraint satisfaction. By introducing the first policy-finetuning based framework for O2O safe RL, which is compatible with many offline and online safe RL methods, our work has the great potential to advance the field towards more efficient and practical safe RL solutions. We provide the official code for Marvel on Github.
@@ 1 Introduction
Safe reinforcement learning (safe RL) [1, 2], prioritizes not only the maximization of rewards but also the adherence to specific safety constraints, enhancing its applicability in real-world scenarios. For example, an autonomous vehicle must reach its destination without exceeding a preset fuel limit. However, solving safe online RL from scratch in fields such as robotics [3, 4], and healthcare [5, 6] is often prohibitive, due to substantial risks and costs caused by the need for extensive interactions with the environment. To address this, offline safe RL [7, 8, 9] has been introduced, enabling the derivation of safe policies from a static dataset [10] without the need for real-time environmental interaction. Nonetheless, offline safe RL faces its own set of limitations: it typically shows limited performance [11], heavily relies on the quality of the offline dataset, and suffers from the impact of out-of-distribution (OOD) actions, restricting its effectiveness across varying scenarios.
@@
The pretraining-and-finetuning paradigm is a well-established strategy in the fields of computer vision and natural language processing, for enabling fast and sample-efficient online learning based on offline pretrained models, particularly with the recent advances in large language models. Following a similar line, offline-to-online reinforcement learning (O2O RL) in the unconstrained case [12, 13, 14] and imitation learning [15, 16] has recently gained prominence. These approaches utilize policies (including Q-functions) derived from offline RL or offline imitation learning, along with offline datasets, to expedite the process of online finetuning. This strategy effectively avoids the extensive environmental interactions required in training policies from scratch. Thus motivated, a key insight is that leveraging the pretraining-and-finetuning paradigm can also potentially facilitate more efficient and practical online safe RL, which however has not been fully explored in the literature. To fill this gap, we seek to answer the following question:
Can we design an effective offline-to-online approach for safe RL to address the limitations of both online safe RL and offline safe RL, thereby enabling fast online safe policy learning?
@@
Figure 1: The “steps” on the x-axis represent the number of policy gradient updates (i.e., optimizer updates). For each gradient update, the agent interacts with the environment for 3 episodes. This convention is followed in the subsequent figures. In (a), we evaluate these methods in the BallCircle environment from Bullet Safety Gym [17], setting the cost limit to 20. As shown, although “Warm Start” begins with a reasonably good initial policy, it performs poorly and overly conservatively, even worse than “From Scratch” in which the policy and Q-functions are initialized randomly. This result suggests that directly finetuning the pretrained policy and Q-functions may actually hinder online learning. In contrast, “Marvel” achieves impressive results, finding a policy with much higher return in just a few online steps while adhering to the cost limit. In (b), t-SNE visualization of state vectors in the environment, reduced to 2D space. Each point represents a state, with rewards uniformly distributed across the space, while costs are sparse, appearing as isolated points or clusters, reflecting their limited association with states.
@@
However, achieving this is highly nontrivial, and simply applying existing O2O algorithms in conventional RL would not work well here due to unique challenges in safe RL. In Fig. 1 (a), ‘Warm Start’ refers to using the offline pretrained policy and Q-networks directly initialize an online safe RL algorithm. ‘From Scratch’ refers to purely online safe RL training. As illustrated, directly finetuning the offline pretrained policy and Q-functions by using standard online safe RL often results in suboptimal performance and, in some cases, complete training failures.
@@
The reasons behind this phenomenon are as follows: a) Erroneous Q-estimations resulted from objective mismatch and offline cost sparsity. In order to avoid explorations beyond the offline data and reduce the extrapolation errors, offline safe RL algorithms typically introduce additional regularizations in the objective function to push up the cost estimates of OOD actions, e.g., VOCE [18] and CPQ [19], leading to a different overall objective from standard online safe RL. More critically, the majority of state-actions in offline datasets for safe RL usually are safe with zero cost (Fig. 1 (b)), resulting in a pretrained cost Q-function that predicts extremely low cost for most in-distribution (IND) state-actions. By erroneously giving high values for OOD state-actions and low values for IND state-actions, the pretrained cost Q-function will conservatively force the online finetuning to stay in the state-action space similar to offline dataset and be reluctant to explore (e.g., cost of “Warm Start” in Fig. 1). b) Mismatch of Lagrange multipliers. Many online safe RL algorithms [20, 21, 7] solve the constrained optimization problem based on the primal-dual approach, which requires a synchronous updating of Lagrange multipliers. Nonetheless, initial values for these multipliers that are matching with the offline policies cannot be obtained from offline safe RL precisely, such that using traditional dual ascent methods to update the Lagrange multipliers may result in slow learning during the online phase even with accurate estimated Q-functions, ultimately degrading the performance of the learned policy. In this work, we seek to design an effective O2O framework for safe RL by addressing these two challenges above.
The main contribution of this work lies in the development of the warM-stArt safe Reinforcement learning with Value prE-aLignment (Marvel) framework, which includes two key components: Value Pre-alignment (VPA) and Adaptive PID Control (aPID). More specifically, VPA adjusts the pretrained Q-functions by re-evaluating the offline policy before online learning based on the offline data only, so as to align the distribution of estimated Q-values with that of true Q-values under the online learning objective for the offline policy. On one hand, by optimistically estimating rewards and pessimistically estimating costs, VPA promotes active exploration during online learning while maintaining the cost below the limit; on the other hand, the active exploration of high-reward state-actions inevitably increases the risk of exploring high-cost state-actions, amplifying the demand of appropriate Langrange multipliers in online finetuning to penalize the cost violations. To jointly handle this risk and the multiplier mismatch problem, instead of directly finding the best initial multipliers, we take an alternative route by seeking to quickly adapt the multipliers to the right values. Particularly, We introduce aPID, an adaptive PID control mechanism that adjusts the Lagrange multipliers based on cost violations, where PID (Proportional-Integral-Derivative) control is a widely used feedback control technique that combines proportional, integral, and derivative components to minimize errors effectively. This approach can quickly stabilize the online finetuning compared to standard dual ascent-based approaches in online safe RL. Extensive experimental results demonstrate the superior performance of our framework over multiple baseline methods on different benchmarks, i.e., Marvel can quickly find a safe policy with the best reward by using only a few online interactions. To the best of our knowledge, Marvel is the first framework that finetunes pretrained offline policy to facilitate fast online learning for safe RL. More importantly, by only leveraging pretrained offline policy/Q-functions and controlling the Lagrange multipliers update, Marvel is compatible with and ready to plug in a lot of state-of-the-art (SOTA) offline and online safe RL approaches.
@@ 2 Preliminaries
Constrained Markov Decision Process.  We consider a standard constrained Markov Decision Process (CMDP) [22, 23], defined by a tuple $(S,A,T,R,C,\gamma,\eta,c_{th})$. Here $S\subseteq\mathbb{R}^{n}$ represents the state space, $A\subseteq\mathbb{R}^{m}$ denotes the action space, $T:S\times A\times S\to[0,1]$ is the transition probability function, $R:S\times A\to[0,R_{\max}]$ is the reward function, and $C:S\times A\to[0,C_{\max}]$ is the cost function. $\gamma\in[0,1]$ is the discount factor, $\eta$ represents the initial state distribution, and $c_{th}$ is the cost threshold that sets the limit on cumulative costs for the policy. A policy $\pi:S\to\mathcal{P}(A)$ is a mapping from states to a probability distribution over actions, where $\pi(a|s)$ denotes the probability of selecting action $a$ in state $s$. In this work, we consider parameterized policies $\pi_{\theta}$, where $\theta$ denotes the parameters of the policy, typically represented by neural networks in deep RL. Given a policy $\pi$, its cumulative reward under policy $\pi$ is defined as $R(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, where $\tau=(s_{0},a_{0},s_{1},a_{1},\dots)$ is a trajectory induced by policy $\pi$, and the expectation is taken over the distribution of trajectories. Similarly, its cumulative cost is defined as $C(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right]$. The Q-function, for a given policy $\pi$, is defined as the expected cumulative reward starting from a state-action pair $(s,a)$ and thereafter following policy $\pi$: $Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. Similarly, the cost Q-function $Q_{c}^{\pi}(s,a)$ is defined as the expected cumulative cost starting from the same state-action pair $(s,a)$ and thereafter following policy $\pi$: $Q_{c}^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. In the context of CMDP, the goal is to find an optimal policy $\pi^{*}$ that maximizes the cumulative reward $R(\pi)$, subject to the constraint that the cumulative cost $C(\pi)$ does not exceed a predefined threshold $c_{th}$. This can be formulated as the following constrained optimization problem:
@@
To solve this, a common approach is to apply the Lagrangian relaxation method [9], where a Lagrange multiplier $\lambda$ is introduced to enforce the cost constraint. This leads to the following primal-dual optimization formulation:
$\min_{\lambda\geq 0}\max_{\pi}\left[R(\pi)-\lambda(C(\pi)-c_{th})\right]$ (2)
which can be solved by iteratively updating the policy $\pi$ and the Lagrange multiplier $\lambda$. Specifically, $\lambda$ is updated by:
$\lambda_{t+1}=\lambda_{t}+\alpha_{\lambda}(C(\pi_{t})-c_{th})$ (3)
where $\alpha_{\lambda}$ is the learning rate.
Online Safe RL.  Primal-dual based algorithms have shown great effectiveness and superior performance in the literature for online safe RL, which can combine a wide range of online unconstrained RL algorithms with the Lagrange multiplier method to create online safe RL algorithms. Without loss of generality, we consider SAC-lag [9] as the online algorithm, a primal-dual based algorithm that integrates the widely used SAC algorithm [24] with the Lagrange multiplier method. More specifically, SAC minimizes the following objectives for the actor (policy) and the critic (Q-function), respectively:
$\mathcal{L}_{\pi}^{SAC}(\theta)=\displaystyle\mathbb{E}_{s\sim d}\displaystyle\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[\alpha\log\pi_{\theta}(a|s)-Q(s,a;\mu)]\$ (4)
$\mathcal{L}_{Q}^{SAC}(\mu)=\displaystyle\mathbb{E}_{(s,a,s^{\prime})\sim d}[(\hat{Q}(s,a;\mu)-y(r,s^{\prime}))^{2}]$ (5)
where $y(r,s^{\prime})=r+\gamma\displaystyle\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}[\hat{Q}(s,a^{\prime};\mu^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})]$, $Q(s,a;\mu)$ is parameterized by $\mu,$ $\hat{Q}(s,a;\mu^{\prime})$ is the target reward Q-function parameterized by $\mu^{\prime}$, $d$ represents the data distribution in the replay buffer, and $\alpha>0$ is some constant. To be applied in online safe RL, SAC-lag adapts SAC by using the Lagrangian method, resulting in the policy optimization objective as follows:
$\mathcal{L}_{\pi}^{SAC}(\theta)=\displaystyle\mathbb{E}_{s\sim d}\displaystyle\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[\alpha\log\pi_{\theta}(a|s)-(Q(s,a)-\lambda{Q_{c}}(s,a))]\$ (6)
@@
Offline Safe RL.  Offline safe RL algorithms typically push up the cost estimations of OOD actions to avoid exploration beyond the offline dataset $\mathcal{D}$. Considering the comprehensive performance across various environments, in this paper we consider the SOTA Lagrangian-based algorithm for offline learning, namely CPQ [19]. More specifically, CPQ first generates OOD actions via a conditional variational autoencoder (CVAE). Then, the cost of the generated OOD actions is increased by minimizing the following loss function for cost critic ($Q_{c}$-function):
@@
$\mathcal{L}_{Q_{c}}^{CPQ}(\mu_{c})=\mathbb{E}_{(s,a,s^{\prime})\sim d}\left[\left(Q_{c}(s,a;\mu_{c})-\left(r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}[\hat{Q}_{c}(s,a^{\prime};\mu_{c}^{\prime})]\right)\right)^{2}\right]-\psi\mathbb{E}_{a\sim d,a\sim\nu}[{Q_{c}}(s,a;\mu_{c})]$
@@
where $Q_{c}(s,a;\mu_{c})$ is parameterized by $\mu_{c},$ $\hat{Q}_{c}(s,a;\mu_{c}^{\prime})$ is the target cost Q-function parameterized by $\mu_{c}^{\prime}$, $\nu$ represents the distribution of OOD actions generated by the CVAE. Additionally, to ensure both constraint safety and in-distribution safety, CPQ updates the reward critic (Q-function) using only state-action pairs that satisfy the cost threshold $l$:
@@
$\mathcal{L}_{Q}^{CPQ}(\mu)=\mathbb{E}_{(s,a,s^{\prime})\sim d}\left[\left(Q(s,a;\mu)-\left(r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}[\mathbb{I}({Q_{c}}(s^{\prime},a^{\prime};\mu_{c})<l)Q(s,a^{\prime};\mu)]\right)\right)^{2}\right]$
where $\mathbb{I}(\cdot)$ is the indicator function, used to filter state-action pairs that satisfy the safety constraints. The policy loss function is given by:
$\mathcal{L}_{\pi}^{CPQ}(\theta)=-\mathbb{E}_{s\sim d}\left[\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[\mathbb{I}({Q_{c}}(s,a;\mu_{c})<l)Q(s,a,\mu)]\right]$ (7)
@@
Similarly, when maximizing the reward, the policy only considers state-action pairs that meet the safety constraints. By assigning a higher cost to OOD actions, CPQ mitigates the OOD problem while meeting safety constraints.
@@
O2O Safe RL.  To the best of our knowledge, Guided Online Distillation [25] is the only work studying O2O safe RL, which leverages a large-scale DT based on GPT-2 [26] as a guide policy to accelerate online learning, by following the idea of Jump-start RL [27]. However, how to achieve fast safe online learning by finetuning a pretrained policy is still not clear. Our work seeks to fill this gap and serves as an initial attempt to spur more interesting studies on policy-finetuning based O2O safe RL without using large models. A more detailed description of related work is delegated to Appendix B.
In this work, our objective is to enable faster and safer policy learning with standard online safe RL methods, by finetuning the policy and Q-functions pretrained using offline safe RL. In principle, any offline safe RL algorithms that output an offline policy and Q-functions can be used here for offline training.
3 Warm-Start Safe RL with Value Pre-Alignment

### v3

# Towards Fast Safe Online Reinforcement Learning via Policy Finetuning

@@ Abstract
High costs and risks involved in extensive environmental interactions hinder the practical application of current online safe reinforcement learning (RL) methods. Inspired by recent successes in offline-to-online (O2O) RL, it is crucial to explore whether offline safe RL can be leveraged to facilitate faster and safer online learning, a direction that has yet to be fully investigated. To fill this gap, we first show that naively applying existing O2O algorithms from standard RL would not work well in safe RL due to two unique challenges: erroneous Q-estimations, resulted from offline-online objective mismatch and offline cost sparsity, and Lagrangian mismatch, resulted from difficulties in aligning Lagrange multipliers between offline and online policies. To address these challenges, we introduce Marvel, the first policy-finetuning based framework for O2O safe RL, comprising two key components that work in concert: Value Pre-Alignment to align the learned Q-functions with the online objective before finetuning, and Adaptive PID Control to effectively adjust the Lagrange multipliers during finetuning. Extensive experiments demonstrate the superior performance of Marvel over related baselines.
@@ 1 Introduction
Safe reinforcement learning (safe RL) (Gu et al., 2022; Garcıa and Fernández, 2015) prioritizes not only reward maximization but also the adherence to safety constraints, enhancing its applicability in real-world scenarios. For instance, an autonomous vehicle must reach its destination without exceeding a fuel limit. However, solving safe online RL from scratch in fields such as robotics (Brunke et al., 2022; Kiran et al., 2021) and healthcare (Yu et al., 2021; Qayyum et al., 2020) is often prohibitive, due to the high risks and costs of extensive environment interactions. To address this, offline safe RL (Zheng et al., 2024; Ray et al., 2019) has been introduced, enabling learning safe policies from a static dataset without online interactions. Yet, offline safe RL faces its own limitations (Ghosh et al., 2022): it typically shows limited performance, relies on the quality of the offline dataset, and suffers from the impact of out-of-distribution (OOD) actions, restricting its effectiveness across varying scenarios.
@@
The pretraining-and-finetuning paradigm is a well-known strategy in the fields of computer vision and natural language processing, for enabling fast online learning based on offline pretrained models. Following a similar line, offline-to-online RL (O2O RL) in the unconstrained setting (Nair et al., 2020; Wang et al., 2024; Zhang et al., 2023b) and imitation learning (Yue et al., 2024; Ross et al., 2011) have recently gained prominence. These approaches utilize policies (including Q-functions) derived from offline learning, along with offline datasets, to expedite online finetuning, which effectively avoids extensive environmental interactions in training policies from scratch. Note that in offline-to-online safe RL, the goal is not large-scale pretraining but rather to obtain a safe warm start from offline data and perform limited online finetuning under safety constraints. Thus motivated, a key insight is leveraging the pretraining-and-finetuning paradigm can also potentially facilitate more efficient and practical online safe RL, which however has not been fully explored in the literature. To fill this gap, we seek to answer the following question: Can we design an effective offline-to-online approach for safe RL to address the limitations of both online and offline safe RL, thereby enabling fast online safe policy learning?
@@
Figure 1: “Steps" on the x-axis represent the number of policy gradient updates (i.e., optimizer updates). For each update, the agent interacts with the environment for 3 episodes. This convention is followed in the subsequent figures. In (a), we evaluate these methods in BallCircle from Bullet Safety Gym (Gronauer, 2022), with a cost limit of 20. As shown, while “Warm Start” begins with a reasonably good initial policy, it performs poorly and overly conservatively, even worse than “From Scratch” where the policy and Q-functions are initialized randomly. This implies that directly finetuning the pretrained policy and Q-functions may hinder online learning. In contrast, “Marvel” achieves impressive results, finding a policy with much higher return in just a few online steps while adhering to the cost limit. In (b), t-SNE visualization of state in the environment, reduced to 2D space. Each point represents a state, with rewards uniformly distributed across the space, while costs are sparse, appearing as isolated points or clusters, reflecting their limited association with states.
@@
However, achieving this is highly nontrivial, and simply applying existing O2O algorithms in conventional RL would not work well here due to unique challenges in safe RL. In Fig. 1 (a), ’Warm Start’ refers to using the offline pretrained policy and Q-networks directly to initialize an online safe RL algorithm. ’From Scratch’ refers to purely online safe RL training. As illustrated, directly finetuning the offline pretrained policy and Q-functions by using standard online safe RL often results in suboptimal performance and, in some cases, complete training failures.
@@
The reasons behind this phenomenon are as follows: a) Erroneous Q-estimations resulting from objective mismatch and offline cost sparsity. To avoid explorations beyond offline data and reduce the extrapolation errors, offline safe RL algorithms typically introduce additional regularizations in the objective function to push up the cost estimates of OOD actions, e.g., VOCE (Guan et al., 2024) and CPQ (Xu et al., 2022), leading to a different overall objective from standard online safe RL. More critically, the majority of state-actions in offline datasets for safe RL usually are safe with zero cost (Fig. 1 (b)), resulting in a pretrained cost Q-function that predicts extremely low cost for most in-distribution (IND) state-actions. By erroneously giving high values for OOD state-actions and low values for IND state-actions, the pretrained cost Q-function will conservatively force the online finetuning to stay in the state-action space similar to offline dataset and be reluctant to explore (e.g., cost of “Warm Start” in Fig. 1). b) Mismatch of Lagrange multipliers. Many online safe RL algorithms (Chow et al., 2018a; Achiam et al., 2017) solve the constrained optimization problem based on the primal-dual approach, which requires a synchronous update of Lagrange multipliers. Nonetheless, initial values for these multipliers that match the offline policies cannot be obtained from offline safe RL precisely. Using traditional dual ascent methods to update the Lagrange multipliers may result in slow online learning even with accurately estimated Q-functions, ultimately degrading the performance of the learned policy. In this work, we seek to design an effective O2O framework for safe RL by addressing these two challenges above.
The main contribution of this work is the warM-stArt safe Reinforcement learning with Value prE-aLignment (Marvel) framework, consisting of two key components: Value Pre-Alignment (VPA) and Adaptive PID Control (aPID). VPA re-aligns pretrained Q-functions to closely match true Q-values of the offline policy under the online learning objective by re-evaluating the offline policy based on offline data alone. To further compensate VPA and address the multiplier mismatch problem, aPID quickly adapts Lagrange multipliers based on cost violations instead of directly finding the best initial multipliers. Experimental results demonstrate Marvel’s ability to quickly learn a safe, high-reward policy with minimal online interactions, marking it as a pioneering solution in O2O safe RL that integrates seamlessly with existing offline and online safe RL methods.
@@ 2 Preliminaries
Constrained Markov Decision Process. We consider a standard constrained Markov Decision Process (CMDP) (Altman, 2021), defined by a tuple $(S,A,T,R,C,\gamma,\eta,c_{th})$. Here $S$ is the state space, $A$ is the action space, $T:S\times A\times S\to[0,1]$ is the transition probability function, $R:S\times A\to[0,R_{\max}]$ is the reward function, and $C:S\times A\to[0,C_{\max}]$ is the cost function. $\gamma\in[0,1)$ is the discount factor, $\eta$ represents the initial state distribution, and $c_{th}$ is the cost threshold that sets the limit on cumulative costs for the policy. A policy $\pi:S\to\mathcal{P}(A)$ is a mapping from states to a probability distribution over actions, where $\pi(a|s)$ denotes the probability of selecting action $a$ in state $s$. Let the policy $\pi_{\theta}$ be parameterized by $\theta$. Given a policy $\pi$, its cumulative reward is defined as $R(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\right]$, where $\tau=(s_{0},a_{0},s_{1},a_{1},\dots)$ is a trajectory induced by policy $\pi$, and the expectation is taken over the distribution of trajectories. Similarly, its cumulative cost is defined as $C(\pi)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\right]$. The reward Q-function for a policy $\pi$ is defined as: $Q^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. Similarly, the cost Q-function $Q_{c}^{\pi}(s,a)$ is defined as the expected cumulative cost starting from the same state-action pair $(s,a)$ and thereafter following policy $\pi$: $Q_{c}^{\pi}(s,a)=\mathbb{E}_{\tau\sim\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}c(s_{t},a_{t})\mid s_{0}=s,a_{0}=a\right]$. In CMDP, the goal is to find an optimal policy $\pi^{*}$ that maximizes $R(\pi)$, subject to the constraint that the cumulative cost $C(\pi)$ does not exceed a predefined threshold $c_{th}$. This can be formulated as the following:
@@
To solve this, a common approach is to apply the Lagrangian relaxation method (Ray et al., 2019) by introducing a Lagrange multiplier $\lambda$ to enforce the cost constraint, leading to the following primal-dual formulation:
$\min_{\lambda\geq 0}\max_{\pi}\left[R(\pi)-\lambda(C(\pi)-c_{th})\right],$ (2)
which can be solved by iteratively updating policy $\pi$ and $\lambda$. $\lambda$ is updated with the learning rate $\alpha_{\lambda}$:
$\lambda_{t+1}=\lambda_{t}+\alpha_{\lambda}(C(\pi_{t})-c_{th}).$ (3)
Online Safe RL.  Primal-dual based algorithms have shown great effectiveness and superior performance in the literature for online safe RL. Without loss of generality, we consider SAC-lag (Ray et al., 2019) as the online algorithm, which integrates the widely used SAC algorithm (Haarnoja et al., 2018) with the Lagrange multiplier method. More specifically, SAC minimizes the following objectives for the actor (policy) and the critic (Q-function), respectively:
$\mathcal{L}_{\pi}^{SAC}(\theta)=\underset{s\sim d,\ a\sim\pi_{\theta}(\cdot|s)}{\mathbb{E}}[\alpha\log\pi_{\theta}(a|s)-Q(s,a;\omega)],$ (4)
$\mathcal{L}_{Q}^{SAC}(\omega)=\underset{(s,a,s^{\prime})\sim d}{\mathbb{E}}[(\hat{Q}(s,a;\omega)-y(r,s^{\prime}))^{2}]$ (5)
where $y(r,s^{\prime})=r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\theta}}[\hat{Q}(s,a^{\prime};\omega^{\prime})-\alpha\log\pi(a^{\prime}|s^{\prime})]$, $Q(s,a;\omega)$ is parameterized by $\omega,$ $\hat{Q}(s,a;\omega^{\prime})$ is the target reward Q-function parameterized by $\omega^{\prime}$, $d$ represents the data distribution in the replay buffer, and $\alpha>0$ is some constant. To be applied in online safe RL, SAC-lag adapts SAC by using the Lagrangian method, such that:
$\textstyle\mathcal{L}_{\pi}^{SAC}(\theta)=$ $\mathbb{E}_{s\sim d}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[\alpha\log\pi_{\theta}(a|s)-(Q(s,a)-\lambda Q_{c}(s,a))].$
@@
Offline Safe RL.  Offline safe RL algorithms typically push up the cost estimations of OOD actions to avoid exploration beyond the offline dataset $\mathcal{D}$. Considering the comprehensive performance across various environments, in this paper we consider the SOTA Lagrangian-based algorithm for offline learning, namely CPQ (Xu et al., 2022). More specifically, CPQ first generates OOD actions via a conditional variational autoencoder (CVAE). The cost of the generated OOD actions is increased by minimizing the following loss function for cost critic ($Q_{c}$-function):
@@
$\textstyle\mathcal{L}_{Q_{c}}^{CPQ}(\omega_{c})=\mathbb{E}_{D}\Big[\left(Q_{c}(s,a;\omega_{c})-\left(c+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}\right.\right.[\hat{Q}_{c}(s^{\prime},a^{\prime};\omega_{c}^{\prime})]))^{2}\Big]-\psi\mathbb{E}_{s\sim D,a\sim\nu}[{Q_{c}}(s,a;\omega_{c})]$
@@
where $D$ represents distribution of $(s,a,s^{\prime})$ in the offline dataset. $Q_{c}(s,a;\omega_{c})$ is parameterized by $\omega_{c},$ $\hat{Q}_{c}(s,a;\omega_{c}^{\prime})$ is the target cost Q-function parameterized by $\omega_{c}^{\prime}$, $\nu$ represents the distribution of OOD actions generated by the CVAE. Additionally, to ensure both constraint safety and in-distribution safety, CPQ updates the reward Q-function using only state-action pairs that satisfy the cost threshold $l$:
@@
$\textstyle\mathcal{L}_{Q}^{CPQ}(\omega)=\mathbb{E}_{D}\big[(Q(s,a;\omega)-(r+\gamma\mathbb{E}_{a^{\prime}\sim\pi_{\theta}(\cdot|s^{\prime})}[\mathbb{I}({Q_{c}}(s^{\prime},a^{\prime};\omega_{c})<l)Q(s^{\prime},a^{\prime};\omega)]))^{2}\big]$
@@
where $\mathbb{I}(\cdot)$ is the indicator function. The policy loss function is defined as: $\mathcal{L}_{\pi}^{CPQ}(\theta)=-\mathbb{E}_{D}\left[\mathbb{E}_{a\sim\pi_{\theta}(\cdot|s)}[\mathbb{I}({Q_{c}}(s,a;\omega_{c})<l)Q(s,a,\omega)]\right]$. Similarly, when maximizing the reward, the policy only considers state-action pairs that meet the safety constraints. By assigning a higher cost to OOD actions, CPQ mitigates the OOD problem while meeting safety constraints.
@@
O2O Safe RL.  To the best of our knowledge, Guided Online Distillation (Li et al., 2024) is the only work studying O2O safe RL, which leverages a large-scale DT based on GPT-2 (Radford et al., 2019) as a guide policy to accelerate online learning, by following the idea of Jump-start RL (Uchendu et al., 2023). However, how to achieve fast safe online learning by finetuning a pretrained policy is still not clear due to the mismatch of Lagrange multipliers and erroneous Q-estimations from objective mismatch and offline cost sparsity.
In this work, we seek to enable faster and safer policy learning by finetuning policies and Q-functions pretrained via offline safe RL, using standard online safe RL methods. Our approach does not rely on large models and is intended as an initial step toward promoting further research on policy-finetuning-based O2O safe RL. In principle, any offline safe RL algorithm that outputs a policy and Q-functions can be used in the offline pretraining stage. More details of related work are provided in Section B.
3 O2O Safe RL with Value Pre-Alignment